OpenStudy (anonymous):

how do you simplify the radical 9 sqrt(1/3

OpenStudy (anonymous):

The only 'simplification' you can do is to move the radical from the denominator by multiplying top and bottom by \(\sqrt{3}\)

OpenStudy (anonymous):

hey jhonny9 can you help me witha problem

jhonyy9 (jhonyy9):

- if you have thought hence : radical9 (1/3)*2 =3/3*2 =1/3

jhonyy9 (jhonyy9):

write it on the question list

OpenStudy (anonymous):

I believe the problem here is: Simplify: \[9\sqrt{\frac{1}{3}}\]

OpenStudy (anonymous):

Which is \[\frac{9}{\sqrt{3}} = \frac{9\sqrt{3}}{3}\]

OpenStudy (anonymous):

Which is \(3\sqrt{3}\)

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

alright i did

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Just remember if you have a radical in the denominator you can clear it by multiplying top and bottom by that same radical. Then you might be able to simplify it.

OpenStudy (anonymous):

This is what I have been doing previously -_______- : 9sqrt 1 / 9sqrt 3 and multipled radical three top and bottom from there. O_O